# k-LLMmeans: Scalable, Stable, and Interpretable Text Clustering via LLM-based Centroids

### v2

@@ Abstract
We introduce k-LLMmeans, a novel modification of the k-means algorithm for text clustering that leverages LLM-generated summaries as cluster centroids, capturing semantic nuances often missed by purely numerical averages. This design preserves the core optimization properties of k-means while enhancing semantic interpretability and avoiding the scalability and instability issues typical of modern LLM-based clustering. Unlike existing methods, our approach does not increase LLM usage with dataset size and produces transparent intermediate outputs. We further extend it with a mini-batch variant for efficient, real-time clustering of streaming text. Extensive experiments across multiple datasets, embeddings, and LLMs show that k-LLMmeans consistently outperforms k-means and other traditional baselines and achieves results comparable to state-of-the-art LLM-based clustering—with a fraction of the LLM calls. Finally, we present a case study on sequential text streams and introduce a new benchmark dataset constructed from StackExchange to evaluate text-stream clustering methods.
@@ 1 Introduction
Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval [37, 42]. A common approach involves generating text embeddings [13, 36, 27, 32, 8, 21] for each document, which are then clustered using traditional algorithms [33]. Among these, k-means [26] is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points. Several alternative methods exist for computing centroids [19, 7, 23], yet they often rely on purely numerical calculations. While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships [34].
We propose k-LLMmeans, a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids. Instead of always computing centroids as the mean of embedding vectors (k-means), in spaced iterations we generate centroids based on textual summaries of the cluster contents. At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster (summarization step). LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents [20]. Figure 1 illustrates the k-LLMmeans algorithm with a single summarization step. The example demonstrates how this step enables the algorithm to escape local minima (common in k-means) and converge to the optimal solution.
Unlike k-means, k-LLMmeans generates a concise textual summary for every centroid at each summarization step, exposing the semantic drift of clusters in real time. This built-in transparency simplifies debugging and validation and is especially valuable in sequential or streaming scenarios, where clusters evolve over time. Figure 2 illustrates one centroid through five summarization updates. To capitalize on this property for large-scale or online data, we adapt mini-batch k-means [38] by inserting our summarization step into a mini-batch k-LLMmeans.
Several studies have explored unsupervised text clustering using LLMs [51, 15, 10, 45, 39, 43, 28], demonstrating state-of-the-art performance across various datasets and benchmarks. These approaches consistently outperform traditional clustering algorithms. Yet despite their impressive empirical gains, existing LLM-based clustering pipelines still face two practical obstacles that limit widespread adoption. First, lack of scalability. Most methods rely on fine-tuning [51], semi-supervised approaches [45], or extensive LLM queries that grow at least linearly with sample size [15, 10], quickly becoming prohibitive for real-world applications. Second, unclear optimization landscape. Most methods braid together bespoke heuristics—hard prompts, similarity thresholds, or greedy merges—without a single, well-defined loss function, leaving convergence behaviour opaque and theoretical guarantees out of reach. We address these two shortcomings with k-LLMmeans: First, by summarizing each cluster in batches we cap LLM usage independent of dataset size; and second, by inheriting the classic k-means objective we retain a provable descent property and a transparent optimization landscape. Regardless of summarization quality, the algorithm will converge to a local minimum.
Our approach is not intended to outperform complex state-of-the-art LLM-based clustering methods, but rather to provide a robust, scalable and LLM-efficient enhancement to the well-established k-means algorithm, preserving its core behavior while using significantly fewer LLM calls than contemporary LLM-based clustering approaches. Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role. Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact.
Figure 1: Illustration of a single-summarization k-LLMmeans algorithm. The first panel displays the text embeddings projected onto the first two principal components along with sample texts, with stars indicating initial random centroids. The second panel shows the local minimum reached by standard k-means. The third panel illustrates the summarization step, where centroids are updated based on LLM-generated summaries of the clustered texts. The final panel presents the resulting clusters after the final k-means step, showing convergence to a more optimal k-LLMmeans solution.
In summary our contributions are as follows:
  • In Section 3 we introduce k-LLMmeans, which enhances the classical k-means algorithm with LLM-powered centroid updates for robust text clustering; Section 4 then builds on this foundation with a mini-batch variant that brings sequential, scalable, and intrinsically interpretable clustering to streaming text data.
  • Section 5 presents extensive simulations demonstrating that k-LLMmeans consistently surpasses k-means and other traditional clustering baselines across a wide range of embeddings and LLMs, while obtaining comparable results to recent LLM-based clustering methods with substantially fewer LLM calls.
  • Section 6 shows that mini-batch k-LLMmeans decisively outperforms streaming variants of k-means, and the case study in Section 7 illustrates how its evolving centroids provide clear, human-readable insight into real-time text streams.
As a by-product of our experiments, we compile a dataset from StackExchange [41] suitable for benchmarking text-streaming analysis methods (Appendix A.2).
Figure 2: An illustration of the dynamic evolution of an LLM-generated centroid during the k-LLMmeans clustering process.
@@ 2 Preliminaries: k-means for text clustering
We can formalizae the text clustering problem. Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
@@
where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $[C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:

### v3

# Summaries as Centroids for Interpretable and Scalable Text Clustering

@@ Abstract
We introduce k-NLPmeans and k-LLMmeans, text-clustering variants of k-means that periodically replace numeric centroids with textual summaries. The key idea—summary-as-centroid—retains k-means assignments in embedding space while producing human-readable, auditable cluster prototypes. The method is LLM-optional: k-NLPmeans uses lightweight, deterministic summarizers, enabling offline, low-cost, and stable operation; k-LLMmeans is a drop-in upgrade that uses an LLM for summaries under a fixed per-iteration budget whose cost does not grow with dataset size. We also present a mini-batch extension for real-time clustering of streaming text. Across diverse datasets, embedding models, and summarization strategies, our approach consistently outperforms classical baselines and approaches the accuracy of recent LLM-based clustering—without extensive LLM calls. Finally, we provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
@@ 1 Introduction
Text clustering is a core problem in natural language processing (NLP), with applications in document organization, topic exploration, and information retrieval [55, 61]. A standard pipeline embeds documents into vectors [17, 54, 38, 46, 10, 29] and then groups them with a clustering algorithm [47]. Among these algorithms, k-means [36] remains ubiquitous, iteratively updating each centroid as the mean of its assigned points. While effective, this purely numerical averaging can blur contextual nuance present in the original texts [52]. Prior work has explored alternative centroid definitions and related objectives [27, 9, 31], yet these approaches remain anchored in vector space, which limits interpretability and can induce semantic drift between centroids and their underlying documents. This motivates centroids that are explicitly textual, aligning prototypes with human-interpretable summaries.
Our proposal.
We introduce a simple modification to k-means: periodically replace numerical centroid updates with summarization steps. Rather than averaging embeddings at every iteration, we compute, at spaced iterations, a textual prototype that summarizes each cluster and then re-embed it with the same encoder to obtain the centroid used for assignments. These summary-as-centroid updates remain inside the standard k-means loop but capture richer contextual meaning. Alternating numerical and summary-based updates yields clusters that are more interpretable and often more semantically coherent. Figure 1 illustrates our proposal and shows how even a single summarization step can redirect k-means toward a qualitatively improved solution.
Figure 1: Illustration of k-NLPmeans/k-LLMmeans with a single summarization step. First panel shows the text embeddings with stars marking the initial centroids; second shows the partition reached after k-means iterations (a local minimum); third performs the summarization step, each previous cluster is summarized into a textual prototype and re-embedded; final panel runs one more k-means iteration using these summaries as centroids, yielding a qualitatively improved partition.
Summarization step.
We consider two families of summarizers: (1) Classical NLP—centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14]—which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans). In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer. While methods from (1) and (2) are standard in summarization, our periodic use of summary-derived centroids within the k-means loop is, to our knowledge, novel and central to the method’s interpretability.
Interpretability and scalability.
Each summarization step produces a concise, human-readable centroid that exposes evolving cluster semantics, simplifying debugging, validation, and labeling. The summarizer is modular: it can be LLM-free (k-NLPmeans) or LLM-assisted under a fixed per-iteration budget that does not grow with dataset size (k-LLMmeans). For large-scale and streaming settings, we adapt mini-batch k-means [56] by inserting summarization steps into the mini-batch update rule, enabling efficient online clustering while retaining interpretability.
Relation to existing LLM-based clustering.
Recent work shows that LLMs can deliver strong unsupervised clustering [73, 20, 13, 64, 58, 62, 39]. However, many pipelines face two practical issues: (i) scalability, as they may rely on semi-supervision, iterative labeling, or a number of LLM calls that grows with dataset size; and (ii) opaque optimization, since they often combine prompts, greedy merges, and similarity thresholds without an explicit objective, making convergence behavior hard to analyze. Our k-LLMmeans addresses both: by summarizing per cluster at spaced iterations, we cap LLM usage independently of dataset size, and by keeping assignments and numeric updates in embedding space, we preserve the standard k-means objective between summary steps. In both k-NLPmeans and k-LLMmeans, if the summaries are weak, the procedure gracefully degrades to vanilla k-means with a suboptimal initialization, retaining its standard local-convergence behavior. Note that our method is LLM-optional: k-NLPmeans replaces the summarizer with classical NLP techniques, incurring no LLM usage.
Contributions.
In summary, we (i) propose k-NLPmeans and k-LLMmeans, a summary-as-centroid variant of k-means that periodically replaces numeric centroids with textual prototypes re-embedded by the same encoder—an LLM-optional design that, by construction, preserves the standard k-means objective between summary steps; interpretability follows from human-readable centroids and transparent intermediate outputs; (ii) extend the approach to mini-batch k-means for efficient, online clustering of streams; (iii) present a comprehensive empirical study across datasets, embeddings, and summarizers, showing consistent gains over classical baselines and competitiveness with recent LLM-based clustering under a dataset-size–independent LLM budget; and (iv) provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
@@ 2 Preliminaries: k-means for text clustering
Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
@@
where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings: